ganeshie8 (ganeshie8):

show that \(a^n+b^n=c^n\) has no solutions in prime numbers for all integers \(n \gt 2\) ( a, b, c can be any prime numbers n is an integer > 2 )

OpenStudy (mathmath333):

is this fermat's last thoerm

ganeshie8 (ganeshie8):

related to that, but we can prove this with some clever elementary algebra.. give it a try :)

OpenStudy (anonymous):

ok, i was thinking u mean a,b,c primes ?

OpenStudy (anonymous):

if yes then i'll assume "c" is odd prime , so "a"need to be odd, and "b" even (or inverse) which implies at least of on them even (composite)

ganeshie8 (ganeshie8):

just to get ourselves acquainted with the equation, maybe first try proving below equation has no solutions in prime numbers : \[a^3 + b^3 = c^3\]

ganeshie8 (ganeshie8):

2 is the only even prime

OpenStudy (anonymous):

oh i thought ur condition a,b,c>2

ganeshie8 (ganeshie8):

a, b, c can be any prime numbers n is an integer > 2

OpenStudy (anonymous):

so , when c is even prime then 2^n=a^n+b^n so we have 2 cases a,b both even or both odd but note 2 is the least prime so ay case would lead to a^n+b^n>2^n

OpenStudy (anonymous):

i'll rewrite in neat way

ganeshie8 (ganeshie8):

you're on right track!

OpenStudy (anonymous):

case 1 :- \(c\) is even prime , \(c=2\) \(a^n+b^2=2^n \rightarrow \) but \(2\) is the least prime \(2^n < a^n+b^n\) case 2:- \(c \) is odd prime , \(c>2\) W.L.O.G \(a \) or \(b\) need to be even and the other is odd let \(a \) be even and \(b\) be odd \(2^n=c^n-b^n\) we wanna show \(c^n-b^n > 2^n \)

ganeshie8 (ganeshie8):

that works!

OpenStudy (cheesecakekitten):

An Honorary Professor of Mathematics is asking US a question?

OpenStudy (cheesecakekitten):

Or are you teaching

OpenStudy (anonymous):

now i'll show induction xD \(c\) and \(b\) are both odd so \(c-b \ge 2\) so for \(n=3\) \(c^3-b^3>2^3\) from ur hint now lets assume its true for k \(c^k-b^2>2^k\) show for k+1 \( c^{k+1}-b^{k+1} > 2^{k+1}\text{note that c is the largest btw them} \) \(\begin{align*} c^{k+1}-b^{k+1}&=c(c^k)-b(b^k) \\ &> c(c^k)-c(b^k) \\ &>c(c^k-b^k) \text{ use induction step} \\ &>c.2^k \text{note c>2} \\ &>2^{k+1} \end{align*}\)

OpenStudy (anonymous):

Tadaaa

OpenStudy (anonymous):

so this work when u prove \(a^3+b^3=c^3\) has no solution i assumed it worked , but if u wanna we can go through it maybe

ganeshie8 (ganeshie8):

Excellent! looks neat.

OpenStudy (anonymous):

\(\text{for } a^3+b^3=c^3 \text{ we only need to show} \\c^3-b^3>2^3\\\begin{align*} c^3-b^3 &= (c-b)(b^2+bc+c^2)\\ &>2.2^2 \\ &>2^3 \end{align*} \)

ganeshie8 (ganeshie8):

if we assume \(c\ne b \gt 2\) we have : \[\begin{align}c^n - b^n &= (c-b)(c^{n-1} + c^{n-2}b + \cdots + b^{n-1}) \\~\\ &\gt 2(c^{n-1} + c^{n-2}b + \cdots + b^{n-1}) \\~\\ &\gt 2(2^{n-1}) \\~\\ &= 2^n \end{align} \]

ganeshie8 (ganeshie8):

that proves that there are no solutions when atleast one of a,b,c is 2

OpenStudy (anonymous):

hehe i was looking for this expansion >.< i

OpenStudy (anonymous):

yes yes ! short and neat

ganeshie8 (ganeshie8):

lol @CheesecakeKitten i just thought this is an interesting problem because it refers to Fermat's last theorem look it up in google, you will find fascinating stories im sure :)

OpenStudy (anonymous):

Fermat was for integers right ?

OpenStudy (cheesecakekitten):

I'd rather not read about math for fun, thank you.

ganeshie8 (ganeshie8):

yes http://mathworld.wolfram.com/FermatsLastTheorem.html

OpenStudy (cheesecakekitten):

WAY too complicated

OpenStudy (cheesecakekitten):

srry!

OpenStudy (cheesecakekitten):

and i'm kinda supposed to be doing my school work

OpenStudy (anonymous):

@CheesecakeKitten you dont have , just dont waste ur time darling ;)

OpenStudy (anonymous):

have to*

OpenStudy (cheesecakekitten):

i can't help it, I'm a procrastinator. :P

ganeshie8 (ganeshie8):

lol

OpenStudy (cheesecakekitten):

haha

OpenStudy (anonymous):

-.-

OpenStudy (cheesecakekitten):

omfg i'm being spammed with messages!!!

OpenStudy (anonymous):

i spammed u :O

OpenStudy (anonymous):

hehe will go now cya ;)

OpenStudy (cheesecakekitten):

lol funny pic

OpenStudy (anonymous):

-.- yeah

ganeshie8 (ganeshie8):

this is my fav

OpenStudy (cheesecakekitten):

XD

ganeshie8 (ganeshie8):

thats hilarious haha (will type the rest later ;) )

OpenStudy (cheesecakekitten):

XD

OpenStudy (cheesecakekitten):

my profile pic. c:

OpenStudy (cheesecakekitten):

just changed it to this an hour or so ago.

ganeshie8 (ganeshie8):

https://www.youtube.com/watch?v=LOMbySJTKpg

OpenStudy (mathmath333):

cool

OpenStudy (aaronandyson):

can you help me @ganeshie8